# Deterministic Bethe state preparation

### v1

@@
We present a quantum circuit that prepares an arbitrary $U(1)$-invariant state on a quantum computer, including the exact eigenstates of the spin-1/2 XXZ quantum spin chain with either open or closed boundary conditions. The algorithm is deterministic, does not require ancillary qubits, and does not require QR decompositions. The circuit prepares such an $L$-qubit state with $M$ down-spins using $\binom{L}{M}-1$ multi-controlled rotation gates and $2M(L-M)$ CNOT-gates.
@@ 1 Introduction
After [19], an algorithm for preparing Bethe states of the spin-1/2 closed XXZ chain based on coordinate Bethe ansatz was formulated [20]. This algorithm, which was later generalized to the open chain [21], was restricted to real Bethe roots, required ancillary qubits, and was probabilistic. Moreover, the success probability was shown to tend exponentially to zero for cases of physical interest [22]. A different approach, based instead on algebraic Bethe ansatz, was subsequently developed for the closed XXZ chain [23, 24]. The latter algorithm is not limited to real Bethe roots, and is deterministic; however, it requires performing QR decompositions of matrices whose size scales exponentially with the number of down-spins.
We present here a new algorithm for preparing arbitrary $U(1)$-invariant $L$-qubit states, including exact Bethe states of the spin-1/2 XXZ chain with either open or closed boundary conditions. Like the algorithms [20, 21], it is based on coordinate Bethe ansatz; however, it is not limited to real Bethe roots, does not require ancillary qubits, and it is deterministic. Moreover, unlike [23, 24], the algorithm is straightforward and does not require QR decompositions. For a state with $M$ down-spins in a chain of length $L$, the circuit uses only $L$ qubits, and has size and depth $\mathcal{O}\left(\binom{L}{M}\right)$. Our algorithm is inspired by an efficient recursive algorithm[25, 26, 27, 28] for preparing Dicke states [29]. Code in Qiskit [30] implementing this algorithm is available as Supplementary Material.
@@ 2 Recursion
That is, we are restricting the sum in (2.4) to permutations $w$ of the form $w=ab$, where $ab$ denotes the concatenation of the strings $a$ and $b$, so $|w\rangle=|a\rangle\otimes|b\rangle$. Note that when $b=\{\}$ is the empty string, $F(b)$ reduces to the normalization given by (2.2). For example, for the state $|B^{4}_{2}\rangle$ (2.3) with $b=0$, we have
@@
$|\psi(0)\rangle=\frac{G(00)}{F(00)}\,f(1100)\,|1100\rangle+\frac{G(10)}{F(10)}\left(f(0110)\,|0110\rangle+f(1010)\,|1010\rangle\right)\,.$ (2.9)
@@ 3 Algorithm
It thus remains to construct a circuit $W_{m}$ that performs (3.5) for all $l$ and $b$. This can be achieved using an operator $\textup{I}_{m,l}$ independent of $b$ with the following properties
@@
for all $b$, where $W_{m}\,|0\rangle^{\otimes(m-l)}\,|1\rangle^{\otimes l}\,|b\rangle_{L-m}$ in (3.8) is given by (3.5). Indeed, it readily follows from these properties that
@@
$\mathcal{U}(m,l)=\prod_{b\in P(L-m,M-l)}u(m,l,b)\,$ (3.12)
whose angles $\theta,\phi,\lambda$ depend on $b$ as follows:
Each $u$-gate has controls on the wires $L-m+l$ and $L-m+l-1$ (the latter of which is present if and only if $l>1$).444 We follow Qiskit conventions, where e.g. $|110\rangle$ corresponds to $|0\rangle$ on wire 00 and $|1\rangle$ on wires $1$ and $2$. Additionally, because the angles depend on $b$, additional controls are placed on wires 00 to $L-m-1$ at the locations where the string $b$ takes the value $1$, to differentiate between different $b$ (this method has been used before to generate states with many free parameters, see e.g. [31]). For example, for $L=4,\,M=2$, $u(m=2,l=1,b=01)$ has controls on the wires $3$ and 00, whereas $u(m=2,l=1,b=10)$ has controls on the wires $3$ and $1$. For the case in which $b$ in (3.12) is unique, these additional controls can be omitted. Complete circuit diagrams for examples with $L=4,\,M=2$ and $L=6,\,M=3$ are displayed in Figs. 2 and 3 in Appendix B. An implementation of this algorithm using Qiskit [30] is available as Supplementary Material.
@@
The number of CNOT-gates in $U_{L}$ is given by
$\sum_{m=2}^{L}\,\,\sum_{l=\max(M+m-L,1)}^{\min(m-1,M)}2\,=2M(L-M)\,,$ (3.15)
as follows from Eqs. (3.7) and (3.11). The number of $u$-gates in $U_{L}$, considering (3.12), is given by
@@
For the special case that all $f(w)=1$, corresponding to Dicke states, our circuit is similar to, but does not reduce to, the one in [25, 26, 27, 28]. This is because the rotation angles in $u(m,l,b)$ are the same for all $b$, so only $M(L-M)$ rotations are needed, and the additional controls on the wires 00 to $L-m-1$ are unnecessary.
@@ 4 Discussion
We have formulated an algorithm for exactly preparing on a quantum computer a $U(1)$-invariant $L$-qubit state consisting of an arbitrary superposition of computational basis states with $M$ ones (down-spins) and $L-M$ zeros (up-spins) (2.1). The circuit is ideally suited for models solved by Bethe ansatz, which have available general recipes for the values $f(w)$ corresponding to the models’ eigenstates for arbitrary values of $L$ and $M$ in terms of Bethe roots $\{k_{1},...,k_{M}\}$. (For eigenstates of non-integrable models, the $f(w)$ values are generally not known.) The circuit size is of the same order as the number of free parameters $f(w)$, suggesting that the circuit is optimal for given values of $L$ and $M$. Nevertheless, for $M=L/2$, the circuit size and depth scale exponentially with $L$. As noted in the Introduction, these states could be used to compute correlation functions [17, 18].
@@ Acknowledgments
We thank Eric Ragoucy and Michelle Wachs for helpful correspondence. RN is supported in part by the National Science Foundation under grant PHY 2310594, and by a Cooper fellowship.
@@ A.2 Open XXZ chain
$\mathcal{H}=-\tfrac{1}{2}\sum_{n=1}^{L-1}\left(\sigma^{x}_{n}\sigma^{x}_{n+1}+\sigma^{y}_{n}\sigma^{y}_{n+1}+\Delta\left(\sigma^{z}_{n}\sigma^{z}_{n+1}-\mathbb{I}\right)\right)-\tfrac{1}{2}\left(h\,\sigma^{z}_{1}+h^{\prime}\,\sigma^{z}_{L}\right)\,,$ (A.8)
@@ Appendix B Circuit diagrams
Figure 2: The complete circuit diagram for the open XXZ Bethe state with $\Delta=0.5,\,h=0.1,\,h^{\prime}=0.3$, see (A.8), for the case $L=4,\,M=2$, and Bethe roots 0.682741, 1.385610.6827411.385610.682741,\,1.385610.682741 , 1.38561. Here the barriers separate the different $W_{m}$, see (3.7) and (3.11). The fourth and fifth $u$-gates correspond respectively to $b=10$ and $b=01$.

### v2

@@
We present an explicit quantum circuit that prepares an arbitrary $U(1)$-eigenstate on a quantum computer, including the exact eigenstates of the spin-1/2 XXZ quantum spin chain with either open or closed boundary conditions. The algorithm is deterministic, does not require ancillary qubits, and does not require QR decompositions. The circuit prepares such an $L$-qubit state with $M$ down-spins using $\binom{L}{M}-1$ multi-controlled rotation gates and $2M(L-M)$ CNOT-gates.
@@ 1 Introduction
After [19], an algorithm for preparing Bethe states of the spin-1/2 closed XXZ chain based on coordinate Bethe ansatz was formulated [20]. This algorithm, which was later generalized to the open chain [21], was restricted to real Bethe roots, required ancillary qubits, and was probabilistic. Moreover, the success probability was shown to tend super-exponentially to zero with the number of down-spins [22]. A different approach, based instead on algebraic Bethe ansatz, was subsequently developed for the closed XXZ chain [23]. The latter algorithm was not limited to real Bethe roots, and was deterministic; however, it required performing QR decompositions of matrices whose size scale exponentially with the number of down-spins. Analytic formulae for the unitaries of [23] were proposed in [24].
We present here a new algorithm for preparing arbitrary $L$-qubit $U(1)$-eigenstates, including exact Bethe states of the spin-1/2 XXZ chain with either open or closed boundary conditions. Like the algorithms [20, 21], it is based on coordinate Bethe ansatz; however, it is not limited to real Bethe roots, does not require ancillary qubits, and it is deterministic. Moreover, unlike [23], our algorithm is explicit, and does not make use of QR decompositions. For a state with $M$ down-spins in a chain of length $L$, the circuit uses only $L$ qubits, and has size and depth $\mathcal{O}\left(\binom{L}{M}\right)$. Our algorithm is inspired by an efficient recursive algorithm[25, 26, 27, 28] for preparing Dicke states [29]. Code in Qiskit [30] implementing this algorithm is available as Supplementary Material.
@@ 2 Recursion
That is, we are restricting the sum in (2.4) to permutations $w$ of the form $w=ab$, where $ab$ denotes the concatenation of the strings $a$ and $b$, so $|w\rangle=|a\rangle\otimes|b\rangle$. The state (2.4) is thus the “$b$-tail” of the Bethe state (2.1). Note that when $b=\{\}$ is the empty string, $F(b)$ reduces to the normalization given by (2.2). For example, for the state $|B^{4}_{2}\rangle$ (2.3) with $b=0$, we have
@@
$|\psi(0)\rangle$ $=G(00)\,|\psi(00)\rangle+G(10)\,|\psi(10)\rangle$
$=\frac{G(00)}{F(00)}\,f(1100)\,|1100\rangle+\frac{G(10)}{F(10)}\left(f(0110)\,|0110\rangle+f(1010)\,|1010\rangle\right)\,.$ (2.9)
@@ 3 Algorithm
It thus remains to construct a circuit $W_{m}$ that performs (3.5) for all $l$ and $b$. This can be achieved using an operator $\textup{I}_{m,l}$ independent of $b$ with the following properties:
@@
for all $b$, where $W_{m}\,|0\rangle^{\otimes(m-l)}\,|1\rangle^{\otimes l}\,|b\rangle_{L-m}$ in (3.8) is given by (3.5). Indeed, it readily follows from these properties that $\textup{I}_{m,l}$ performs the mapping (3.5) for fixed $l$, and that
@@
$\mathcal{U}(m,l)=\prod_{b\in P(L-m,M-l)}u(m,l,b)\,$ (3.12)
whose angles $\theta,\phi,\lambda$ depend on $b$ as follows:
Each $u$-gate acts on wire $L-m$, and has controls on the wires $L-m+l$ and $L-m+l-1$ (the latter of which is present if and only if $l>1$).444 We follow Qiskit conventions, where e.g. $|110\rangle$ corresponds to $|0\rangle$ on wire 00 and $|1\rangle$ on wires $1$ and $2$. Additionally, because the angles depend on $b$, additional controls are placed on wires 00 to $L-m-1$ at the locations where the string $b$ takes the value $1$, to differentiate between different $b$ [^5] This method has been used before to generate states with many free parameters, see e.g. [31]. For example, for $L=4,\,M=2$, $u(m=2,l=1,b=01)$ has controls on the wires $3$ and 00, whereas $u(m=2,l=1,b=10)$ has controls on the wires $3$ and $1$. For the case in which $b$ in (3.12) is unique, these additional controls can be omitted. Complete circuit diagrams for examples with $L=4,\,M=2$ and $L=6,\,M=3$ are displayed in Figs. 2 and 3 in Appendix B. An implementation of this algorithm using Qiskit [30] is available as Supplementary Material.
@@
The number of CNOT-gates in $U_{L}$, as follows from Eqs. (3.7) and (3.11), is given by
$\sum_{m=2}^{L}\,\,\sum_{l=\max(M+m-L,1)}^{\min(m-1,M)}2\,=2M(L-M)\,.$ (3.15)
The number of $u$-gates in $U_{L}$, considering (3.12), is given by
@@
For the special case that all $f(w)=1$, corresponding to Dicke states, our circuit is similar to, but does not reduce to, the one in [25, 26, 27, 28]. This is because the rotation angles in $u(m,l,b)$ are the same for all $b$, so only $M(L-M)$ total rotations are needed, and the additional controls on the wires 00 to $L-m-1$ are unnecessary.
@@ 4 Discussion
We have formulated an algorithm for exactly preparing on a quantum computer an $L$-qubit $U(1)$-eigenstate consisting of an arbitrary superposition of computational basis states with $M$ ones (down-spins) and $L-M$ zeros (up-spins) (2.1). The circuit is ideally suited for models solved by Bethe ansatz, which have available general recipes for the values $f(w)$ corresponding to the models’ eigenstates for arbitrary values of $L$ and $M$ in terms of Bethe roots $\{k_{1},...,k_{M}\}$. However, these coefficients $f(w)$, of which there are (${L\choose M}$ many, are each expressed as a sum over $M!$ terms for the closed chain (A.5), and $2^{M}M!$ terms for the open chain (A.12), and are thus expensive to compute classically for large $M$. (For eigenstates of non-integrable models, the $f(w)$ values are generally not known.) The circuit size is of the same order as the number of free parameters $f(w)$, suggesting that the circuit is optimal for given values of $L$ and $M$. Nevertheless, for $M=L/2$, the circuit size and depth scale exponentially with $L$. As noted in the Introduction, these states could be used to compute correlation functions [17, 18].
@@ Acknowledgments
We thank Esperanza Lòpez, Sergei Lukyanov, Eric Ragoucy, Roberto Ruiz, Germàn Sierra, Alejandro Sopena and Michelle Wachs for helpful correspondence or discussions. RN is supported in part by the National Science Foundation under grant PHY 2310594, and by a Cooper fellowship.
@@ A.2 Open XXZ chain
$\mathcal{H}=-\tfrac{1}{2}\sum_{n=1}^{L-1}\left(\sigma^{x}_{n}\sigma^{x}_{n+1}+\sigma^{y}_{n}\sigma^{y}_{n+1}+\Delta\left(\sigma^{z}_{n}\sigma^{z}_{n+1}-\mathbb{I}\right)\right)-\tfrac{1}{2}\left(h\,\sigma^{z}_{1}+h^{\prime}\,\sigma^{z}_{L}\right)+\tfrac{1}{2}(h+h^{\prime})\mathbb{I}\,,$ (A.8)